Janet is 3 years older than her sister Julie. Janet's brother is 8 years younger than their sister Julie. The sum of all their ages is 55. How old are Janet, Julie and their brother?
Answer by TimothyLamb(4379)

x = janet age now
y = julie age now
z = brother age now
---
x + y + z = 55
x = y + 3
z = y - 8
---
put the system of linear equations into standard form
---
x + y + z = 55
x - y = 3
y - z = 8
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = janet age now = 23
y = julie age now = 20
z = brother age now = 12
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php